### v1

# Optimizing quantum battery performance by reducing battery influence in charging dynamics

Abstract
Quantum batteries have emerged as promising devices that work within the quantum regime and provide energy storage and power delivery. We investigate the interplay between the battery and charger Hamiltonians, with a particular focus on minimizing the battery’s influence during the charging dynamics. To achieve this, we introduce a control parameter that allows us to suppress the battery’s contribution during the charging dynamics. We explore various configurations, including a non-interacting many-body battery with an interacting many-body charger, an interacting battery with a non-interacting charger, and systems where both the battery and charger are interacting many-body systems. Our results reveal a notable enhancement in stored energy and charging power when the battery’s influence is suppressed, underscoring the pivotal role of the charger in optimizing performance. Remarkably, across all scenarios, we observe that the presence of the battery’s countereffect within the charger Hamiltonian consistently leads to improved storage characteristics, highlighting a fresh direction in designing efficient quantum batteries.
I Introduction
A quantum battery is a device designed to harness quantum properties to improve the efficiency of energy storage and retrieval. Traditional batteries store energy in chemical reactions [1, 2, 3, 4, 5], while quantum batteries would utilize the principles of quantum mechanics to store and release energy in a more efficient and potentially higher-capacity manner [6, 7, 8, 9, 10, 11, 12, 13]. The basic idea behind a quantum battery is that quantum systems can exist in superposition states, meaning they can be in a combination of multiple states simultaneously. This property could potentially allow quantum batteries to store energy in a more distributed and efficient way compared to traditional batteries. Additionally, quantum entanglement, where particles become correlated in such a way that the state of one particle determines the state of the other particle, might enable quantum batteries to further enhance their energy storage capacity.
Quantum batteries have attracted significant attention in recent years [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. The primary goal of these efforts is to demonstrate how quantum mechanical systems and operations can enhance the stored energy or charging power of a battery. A variety of approaches—ranging from single-qubit protocols to many-body systems—have been explored to showcase this quantum advantage [10, 19, 20, 22, 16, 8, 23]. Notable models include the Dicke quantum battery, which achieves modest speedups due to collective effects [16, 17], and qubit-based schemes that demonstrate an $N$-fold power advantage, pointing to strong scalability [8]. Moreover, coherent quantum charging protocols have shown superior performance compared to classical strategies, with significantly enhanced power and robustness against decoherence and loss [24]. Recent studies have also underscored the importance of interaction topology: optimal performance in fermionic systems is observed when the interaction connectivity aligns with the interaction order [23]. In addition, alternative charging mechanisms such as those based on the Jaynes–Cummings model reveal that non-Gaussian cavity states can minimize energy fluctuations and achieve near-perfect charging fidelity [25]. These theoretical developments have been paralleled by experimental advances, with successful realizations using superconducting circuits and trapped ions [26, 27, 28, 29]. Periodic modulation techniques have also been applied to spin systems for efficient charging [30], and the role of dissipation in open quantum systems has been investigated as a tool for enhancing performance [31, 32, 33, 34, 35, 36, 37, 38, 12]. Notably, global charging methods—where all cells are charged simultaneously—have been shown to offer the strongest quantum advantage, with power scaling quadratically with the system size [21].
There has been a lack of discussion on the combined effect of the battery and charger in the charging dynamics, particularly in the context of spin systems. Addressing this gap, our study explores the interplay between the battery and charger during the charging process, emphasizing their joint influence on energy storage and power output. Our objective is to understand how optimizing their interactions can enhance overall performance while overcoming inherent limitations in the battery’s contribution.
In quantum spin systems, both the battery and the charger play a role in charging dynamics for storing energy and power. However, we observe that the battery’s ability to achieve maximum storage energy and power is often constrained by the charger, which diminishes the battery’s effect during the charging process. To address this, we propose introducing an additional term in the charger Hamiltonian to counteract this limitation. By incorporating the battery’s countereffect into the charger, we enhance both storage energy and power across all considered configurations. This approach shifts the primary responsibility for energy storage to the charger, allowing the battery to adopt a more passive role, thereby optimizing the charging process and maximizing the efficiency of quantum batteries.
The manuscript is organized as follows: Section II introduces the setup involving the combined battery and charger system and defines the key physical quantities used in the analysis. Section III presents the results of energy storage and power calculations. Specifically, Subsection III.1 examines the scenario where the battery is non-interacting and the charger is interacting, while Subsection III.2 addresses the case of an interacting battery coupled to a non-interacting charger. Subsection III.3 explores the dynamics when both the battery and the charger are interacting. Finally, Section IV summarizes the main findings and concludes the manuscript.
II Set-up
A quantum battery typically comprises a finite number of quantum systems, such as spins, that can store and transfer energy at the quantum level. The charger, also a quantum system composed of spins, is designed to inject energy into the battery. The dynamics of these spin-based systems are described by their respective Hamiltonians. The battery Hamiltonian, $\hat{H}_{B}$, describes the internal energy levels and/or interactions within the battery, whereas the charger Hamiltonian, $\hat{H}_{C}^{e}$, encapsulates the dynamics and properties of the charger.
To initiate the charging process, the charger is connected to the battery, allowing them to interact. This interaction results in the formulation of the total charging Hamiltonian, which combines the contributions from both the battery and the charger. It is expressed as the sum of the battery Hamiltonian, $\hat{H}_{B}$, and the charger Hamiltonian, $\hat{H}_{C}^{e}$, capturing the dynamics of energy exchange between the two systems. Mathematically, it is expressed as
$\hat{H}=\hat{H}_{B}+\hat{H}_{C}^{e}.$ (1)
To extend the formulation, we treat the charger Hamiltonian in a unique way by considering both interacting and non-interacting forms. In each case, we incorporate the countereffect of the battery into the charger Hamiltonian to account for the mutual influence between the battery and charger during the charging dynamics. The strength of this countereffect is controlled by a parameter, $\lambda$, referred to as the battery’s countereffect parameter. This modified Hamiltonian is termed the effective charger Hamiltonian, as it accounts for both the intrinsic dynamics of the charger and the influence of the battery. Mathematically, it is given as:
$\hat{H}_{C}^{e}=\hat{H}_{C}-\lambda\hat{H}_{B}.$ (2)
This effective charger Hamiltonian is introduced during the charging process, where it operates alongside the battery Hamiltonian. By incorporating the battery’s countereffect into the charger dynamics, the total Hamiltonian governing the charging process is updated. Substituting the effective charger Hamiltonian into the total Hamiltonian [Eq. 1], we obtain:
$\hat{H}=\hat{H}_{B}+\hat{H}_{C}-\lambda\hat{H}_{B}=\hat{H}_{C}+(1-\lambda)\hat{H}_{B},$ (3)
123NN-145
Figure 1: Illustration of all-to-all interaction of spins.
Physically interesting range of $\lambda$ lies between $[0,1]$. When $\lambda=0$, the battery fully contributes to energy storage and power during the charging process. Conversely, when $\lambda=1$, the battery’s influence is completely nullified during charging. For intermediate values of $\lambda$, the battery’s effect is progressively reduced, with the extent of this reduction determined by the strength of $\lambda$.
To analyze the counter effect of the battery during the charging process, we calculate two key metrics: storage energy and power. The energy stored in the battery during the charging process is defined as:
$\Delta E=\mathrm{Tr}[\hat{\rho}(t)\hat{H}_{B}]-\mathrm{Tr}[\hat{\rho}(0)\hat{H}_{B}],$ (4)
where $\hat{\rho}(0)=|\psi_{0}\rangle\langle\psi_{0}|$, and $|\psi_{0}\rangle$ is the ground state of the battery Hamiltonian. The time evolution of the system’s state is given as: |$\ket{\psi(t)}=\hat{U}(t)\ket{\psi_{0}}$, where $\hat{U}(t)=e^{-i\hat{H}t}$ is the time-evolution operator.
The storage power quantifies the rate at which energy is transferred into the system during the charging process and is given by:
$P=\frac{\Delta E}{T}.$ (5)
Here, $P$ represents the average power over the entire charging duration, and $T$ denotes the total charging time. This formulation provides a measure of how efficiently energy is accumulated, offering insight into the performance and effectiveness of the energy storage process.
III result
To evaluate the performance of the quantum battery, we compute the stored energy and charging power across three distinct scenarios. In the first scenario, the battery is modeled as a non-interacting spin system, while the charger consists of interacting spins. In the second, the battery is an interacting spin system, and the charger is non-interacting. The third configuration involves both the battery and the charger as interacting spin systems. These setups allow us to systematically explore how the battery’s intrinsic properties influence the energy storage and power delivery during the charging process. We begin our analysis with the first scenario.
III.1 Non-interacting battery and interacting charger Hamiltonian
The non-interacting battery is modeled as a collection of spins aligned along the $z$-direction, with each spin having a strength $h$. The battery Hamiltonian is defined as:
$\hat{H}_{B}=h\sum_{j=1}^{N}\hat{\sigma}_{j}^{z},$ (6)
where $N$ is the total number of spins. In its spectral decomposition, the battery Hamiltonian can be expressed as:
$\hat{H}_{B}=h\sum_{n=1}^{2^{N}}E_{n}|E_{n}\rangle\langle E_{n}|,$ (7)
where $E_{n}$ are the eigenvalues, ranging from $-N$ (ground state energy) to $N$ (highest energy level). The total number of energy levels is $2^{N}$, with the ground state and the highest energy state being non-degenerate. The remaining energy levels exhibit degeneracy, given by ${}^{N}C_{l}$, where $l=1,2,\dots,N-1$, corresponding to intermediate energy levels.
The goal of charging the battery is to transition it from the ground energy state |$\ket{E_{1}}$ with energy $-Nh$ to the highest energy state |$\ket{E_{2^{N}}}$ with energy $Nh$. The maximum storage energy $\Delta E_{\rm max}$ is defined as:
[…]
$=\Big{[}Nh-(-Nh)\Big{]}=2Nh.$
The maximum storage energy of the non-interacting battery, as described by Eq. (6), is given by $2Nh$.
Figure 2: Non-interacting battery with ATA interacting Ising spin charger. (a) $\Delta E$ and (b) $P$ vs. $t$ for different values of $\lambda$ with fixed system size $N=10$. (c1) and (c2) show $\Delta E$ vs. $t$ for systems with (c1) odd and (c2) even $N$ with fixed $\lambda=1$. (d) $P$ vs. $t$ for various $N$. The parameters used are $J=h=1$, with periodic boundary conditions applied.
Figure 3: Maximum storage (a) energy and (b) power as functions of the battery’s countereffect strength $\lambda$, for two interaction scenarios and their special cases, with system size fixed at $N=10$. Maximum storage (c) energy and (d) power as functions of system size $N$, with the battery counter strength fixed at $\lambda=1$. In all cases, the parameters used are $J=h=1$, and $\gamma=0.5$.
To examine the influence of the battery’s countereffect on its storage energy and power, we consider an all-to-all (ATA) interacting Ising spin system as an interacting charger, as illustrated in Fig. 1. The interaction strength between a spin and its $k$-th neighbor decays as $\frac{1}{2^{k-1}}$. The corresponding Hamiltonian is given by:
$\hat{H}_{C}=J\sum_{j=1}^{N}\left(\sum_{k=1}^{\mathcal{K}}\frac{1}{2^{k-1}}\sigma_{j}^{x}\sigma_{j+k}^{x}\right),$ (9)
where $\mathcal{K}$ represents the maximum interaction range. We assume periodic boundary conditions, i.e., $\hat{\sigma}_{N+k}=\hat{\sigma}_{k}$, to preserve translational symmetry and eliminate edge effects. To avoid redundancy under periodic boundary conditions, we fix $\mathcal{K}=\frac{N-1}{2}$ for odd $N$ and $\mathcal{K}=\frac{N}{2}$ for even $N$, ensuring that each spin pair is included only once.
We calculate the storage energy of the non-interacting battery Hamiltonian with the ATA connected Ising spin system as the charger Hamiltonian for different values of $\lambda$ ranging from $[0,1]$. It is observed that for all values of $\lambda$ (except at $\lambda=1$), the storage energy increases and eventually oscillates about a particular value with random amplitudes and frequencies. At $\lambda=1$, the storage energy increases and saturates at a particular value. However, at a certain time, a sudden jump occurs, leading to the highest storage energy, $\Delta E_{\rm max}=2hN$. This suggests that at this value of $\lambda$, the battery reaches its highest energy state [Fig. 2(a)].
The storage power exhibits a similar behavior for all values of $\lambda$, increases, reaches its maximum value at a specific time, and then starts to decrease. As the value of $\lambda$ increases, the maximum storage power also increases. [Fig. 2(b)].
The storage energy is analyzed as a function of the system size, with the parameter $\lambda$ fixed at $1$, revealing an odd-even effect in its behaviour. For odd system sizes, the storage energy increases, saturates at a particular value with very small oscillation amplitudes, and then decreases to zero at a specific time. The region of saturation also depends on the system size. As the system size increases, the time regime of constant energy increases. In other words, the storage energy exhibits periodic behavior, with the period of the storage energy increasing as the system size grows [Fig. 2($c_{1}$)]. Additioanlay, maxima of the storage energy also increase with increasing system size. Conversely, for even system sizes, the storage energy increases as time progresses, saturates for a fixed period, and reaches its maximum value at a specific time. The maximum value of the storage energy suggests that, in this scenario, the battery attains its highest energy state, $2hN$. Similar to odd system sizes, the storage energy for even system sizes also exhibits periodic behavior, with the period increasing as the system size grows [ Fig. 2($c_{2}$)].
The storage power of the battery is analyzed as a function of the system size with a fixed value of $\lambda=1$. It is observed that while the storage power depends on the system size, it does not exhibit the odd-even effect. For all system sizes, the storage power increases over time, reaches a maximum value, and then begins to decline. The maximum storage energy increases with the growth of the system size. [Fig. 2(d)].
Optimum storage energy and power
To better understand how the counter effect influences the charging process, we investigate how the maximum storage energy and charging power vary with the counter effect parameter $\lambda$. We also examine how these quantities scale with the system size. For this purpose, we consider two distinct types of charger Hamiltonians.
The first type is defined in Eq. (9) and corresponds to the Ising spin system with ATA interactions. As a special case, we also study the nearest-neighbor (NN) Ising spin system by setting the interaction range $\mathcal{K}=1$. In this case, the charger Hamiltonian simplifies to the standard NN Ising model:
$\hat{H}_{C}=J\sum_{j=1}^{N}\hat{\sigma}_{j}^{x}\hat{\sigma}_{j+1}^{x}.$ (10)
The second type of charger is qualitatively different and is based on the XY spin system with ATA interactions. The Hamiltonian is given by:
$\hat{H}_{C}=$ $(1+\gamma)\sum_{j=1}^{N}\left(\sum_{k=1}^{\mathcal{K}}\frac{1}{2^{k-1}}\hat{\sigma}_{j}^{x}\hat{\sigma}_{j+k}^{x}\right)$ (11)
$+(1-\gamma)\sum_{j=1}^{N}\left(\sum_{k=1}^{\mathcal{K}}\frac{1}{2^{k-1}}\hat{\sigma}_{j}^{y}\hat{\sigma}_{j+k}^{y}\right),$
where $\gamma$ is the anisotropy parameter and $\mathcal{K}$ denotes the interaction range. We also consider the NN case of the Eq. (11)that is obtained by setting $\mathcal{K}=1$, resulting in the following Hamiltonian:
$\hat{H}_{C}=(1+\gamma)\sum_{j=1}^{N}\hat{\sigma}_{j}^{x}\hat{\sigma}_{j+1}^{x}+(1-\gamma)\sum_{j=1}^{N}\hat{\sigma}_{j}^{y}\hat{\sigma}_{j+1}^{y}.$ (12)
By analyzing these different charger configurations, we systematically compare their influence on the battery’s maximum storage energy and power. This comparison provides valuable insights into how the type and range of interactions affect the performance and optimization of quantum batteries.
By employing two types of charger Hamiltonians—Ising and XY models—with both ATA interactions and their special case of NN interactions, we investigate the maximum storage energy and power of the battery as functions of the countereffect strength and the system size.
In all considered cases, the maximum storage energy increases with increasing $\lambda$ and reaches its optimum at $\lambda=1$. [Fig. 3(a)]. For the Ising spin system with ATA interactions, the maximum storage energy reaches the upper bound of $2hN$ at $\lambda=1$. In contrast, when restricted to NN interactions, the storage energy only reaches half this value, i.e., $hN$, at the same $\lambda$. In the case of the XY ATA interacting charger, the battery attains a maximum storage energy greater than $hN$, though not reaching $2hN$. For the XY NN case, the energy remains below $hN$ even at $\lambda=1$. These observations highlight that long-range interactions (especially ATA) significantly enhance the energy storage capacity of the quantum battery. Thus, $\lambda=1$ marks an optimal point where energy transfer during the charging protocol becomes most efficient, particularly in systems with long-range interactions.
The behavior of the storage power also supports the efficiency of the charging protocol with increasing countereffect. As shown in Fig. 3(b), the maximum storage power increases approximately linearly with $\lambda$ in all the considered cases, reaching its optimum value at $\lambda=1$. These results reinforce the conclusion that the countereffect of the battery—modeled by the parameter $\lambda$—facilitates more efficient energy transfer and charging dynamics. In particular, $\lambda=1$ consistently yields the best performance in terms of both storage energy and power across all interacting charger configurations. Among the different interaction types, the Ising spin systems (both NN and ATA) exhibit higher storage power compared to their XY counterparts. This indicates that the nature of the spin-spin interaction affects the storage capacity.
Furthermore, we investigate how the system size $N$ influences the maximum storage energy and power, keeping the countereffect parameter fixed at $\lambda=1$. Interestingly, for chargers with ATA interactions, we observe an odd-even effect in the storage energy behavior. In the Ising ATA case, when the system size $N$ is even, the maximum storage energy reaches its upper bound of $2hN$, whereas for odd $N$, it only attains half of that value, i.e., $hN$. On the other hand, in the NN Ising interaction case, the storage energy consistently reaches half of the maximum possible value, $hN$, regardless of system size. A similar odd-even behavior appears in the XY ATA interacting charger. For odd system sizes, the storage energy remains below half of the maximum value ($<hN$), while for even system sizes, it exceeds half the maximum ($>hN$). In contrast, the XY NN configuration yields a storage energy that remains below $hN$ for all system sizes [Fig. 3(c)].
The storage power exhibits a linear increase with system size $N$ across all the considered charger configurations. Among them, the Ising ATA interaction yields the highest storage power, demonstrating superior charging efficiency. In contrast, the XY ATA configuration results in the lowest storage power, indicating relatively less efficient energy transfer dynamics [Fig. 3(d)].
III.2 Interacting battery and non-interacting charger Hamiltonian
Figure 4: (a) $\Delta E$ and (b) $P$ vs. $t$ for NN Ising interacting Hamiltonian as a battery and non-interacting Hamiltonian as a charger. (c) $\Delta E$ and (d) $P$ vs. $t$ for NN interacting XY spin system as a battery and non-interacting Hamiltonian as a charger. Parameterse are: $N=12$, $J=1$, $h=1$, and $\gamma=0.5$. Periodic boundary conditions are considered.
To generalize our observation that the countereffect of a battery in the charging dynamics enhances storage energy and power, we now consider a reversed setup: an interacting battery coupled to a non-interacting charger. The interacting battery Hamiltonians considered are of two types: Ising spin system with NN interaction defined by Eq.(10), i.e., $\hat{H}_{B}=J\sum_{j=1}^{N}\sigma_{j}^{x}\sigma_{j+1}^{x}$, and XY spin system with NN interaction defined by Eq.(12), i.e., $\hat{H}_{B}=(1+\gamma)\sum_{j=1}^{N}\sigma_{j}^{x}\sigma_{j+1}^{x}+(1-\gamma)\sum_{j=1}^{N}\sigma_{j}^{y}\sigma_{j+1}^{y}$. The non-interacting charger Hamiltonian is given by Eq. (6), i.e., $\hat{H}_{C}=h\sum_{j=1}^{N}\hat{\sigma}_{j}^{z}$.
We analyze the storage energy and power of the battery in two scenarios: one where the battery’s influence on the charger is present ($\lambda=0$) and another where it is absent ($\lambda=1$). The results demonstrate that, for both Ising and XY interacting batteries, storage energy and power are notably higher when $\lambda=1$, underscoring the beneficial role of the battery’s countereffect in the charging dynamics. Additionally, the Ising model consistently yields superior performance over the XY model in both metrics, as shown in Fig. 4(a–d).
III.3 Interacting battery and interacting charger Hamiltonian
We consider the scenario where both the battery and charger are governed by interacting Hamiltonians. First, we analyze the case where the battery is modeled by an Ising spin system with NN interactions [Eq. (10)], and the charger is described by an NN XY-interacting Hamiltonian [Eq. (12)]. We calculate the storage energy and power in two situations: when the battery’s influence is present in the charging dynamics ($\lambda=0$) and when it is absent ($\lambda=1$). Our results indicate that the presence of the counter effect of battery during charging leads to an advantage in both the stored energy and the power [Fig. 5(a,b)].
Next, we examine the reverse configuration: the battery is now represented by an NN XY-interacting spin system [Eq. (12)], while the charger is governed by an NN Ising-type interaction [Eq. (10)]. In this case, we again compute the storage energy and power, and we find that when the battery’s influence is absent in the charging dynamics ($\lambda=1$), the system exhibits improved performance in terms of both energy storage and power [Fig. 5(c,d)].
In nutshell, across all scenarios considered, the presence of the battery’s counter effect in the charging dynamics consistently provides an advantage in both energy storage and power. This advantage is consistent for being independent of the interaction and noniteration of the battery and charger Hamiltonians.
Figure 5: (a) $\Delta E$ and (b) $P$ as a function of time $t$ for a NN interacting Ising spin system as the battery and NN interacting XY spin system as the charger Hamiltonian. (c) $\Delta E$ and (d) $P$ vs. $t$ for a NN interacting XY spin system as the battery and NN interacting Ising spin system as the charger Hamiltonian. The system parameters are $N=12$, $J=1$, and $\gamma=0.5$. Periodic boundary conditions are considered.
IV Conclusion
In this work, we explored the role of the battery’s countereffect in the charging dynamics of quantum batteries composed of spin systems. By analyzing various configurations involving both non-interacting and interacting Hamiltonians for the battery and charger, we demonstrated how their interplay significantly impacts storage energy and power.
We began by examining a non-interacting battery coupled to an interacting charger modeled by an ATA Ising spin system. Our results revealed that incorporating the battery’s countereffect into the dynamics ($\lambda=1$) leads to optimal performance, with the storage energy reaching its theoretical maximum $2Nh$ and the power also peaking. Notably, an odd-even parity effect emerged in energy storage for increasing system sizes, while power remained parity-independent.
To further understand the influence of interaction types and ranges, we compared Ising and XY charger models with both ATA and NN interactions. Across all these cases, the optimum storage energy and power consistently occurred at $\lambda=1$, emphasizing the constructive role of the countereffect in enhancing energy transfer. Both quantities exhibited linear scaling with system size, confirming the scalability of our protocol.
Extending the analysis to interacting batteries with non-interacting chargers, where the battery is governed by either Ising or XY NN interactions, we observed that the inclusion of the countereffect continued to offer superior energy and power outcomes. This advantage persisted even in the most general scenario where both the battery and charger are interacting systems, reinforcing the robustness of our findings across diverse interaction frameworks.
Interestingly, in cases where the charger is modeled with ATA interactions, the charging power peaks beyond the physically meaningful range ($\lambda>1$). Although this shift lies outside the regime of practical relevance, it suggests the presence of underlying physical mechanisms that merit deeper investigation. A detailed discussion of this anomaly is provided in Appendix A.
Altogether, our study establishes the critical importance of retaining the battery’s countereffect in the charger Hamiltonian for optimizing quantum battery performance. These insights offer valuable guidance for the design of efficient and scalable quantum energy storage devices and open new avenues for exploring interaction-driven enhancements in quantum thermodynamic protocols.
Appendix A Charging dynamics beyond physical regime
Figure 6: Storage power $P(t)$ as a function of time $t$ for ATA interacting (a) Ising and (b) $XY$ charger Hamiltonian, shown for various values of the countereffect parameter $\lambda$. The insets display the corresponding maximum storage power as a function of $\lambda$. Parameters used: $J=h=1$, $N=10$, and $\gamma=0.5$.
In the main manuscript, we investigate the charging dynamics of quantum batteries within the physically acceptable range of the countereffect parameter, $\lambda\in[0,1]$. Within this regime, both the storage energy and power attain their maximum at $\lambda=1$. Interestingly, our analysis further reveals that the storage power reaches an even higher value beyond this physical regime ($\lambda>1$) in the case of ATA interacting Ising and $XY$ charger Hamiltonians. However, the storage energy still achieves its maximum at $\lambda=1$. Although the region $\lambda>1$ falls outside the standard physical constraints, we emphasize this result to highlight its potential implications and to encourage future investigations into such unconventional regimes.
We analyze the storage power for a non-interacting battery defined by Eq. (6) and ATA-interacting Ising and XY chargers described by Eq. (9) and Eq. (11), respectively, considering the countereffect parameter of the battery in the range $\lambda\in[0,5]$. In both cases, for all values of $\lambda$, the storage power exhibits a consistent trend: it initially increases, reaches a maximum value, and subsequently decreases after a certain time [Fig. 6(a,b)]. The maximum storage power, $P_{\rm max}$, depends on $\lambda$ and increases with it, attaining its peak at $\lambda=1.6$ for the ATA interacting Ising charger and $\lambda=2.8$ for the ATA interacting XY charger. Beyond these points, the storage power decreases with further increases in $\lambda$ [inset of Fig. 6(a,b)]. Remarkably, the highest $P_{\rm max}$ values occur outside the physically accepted regime ($\lambda>1$).

### v2

@@
-Optimizing quantum battery performance by reducing battery influence in charging dynamics
+System versus charger in performance optimization of quantum batteries
 Abstract
 Quantum batteries have emerged as promising devices that work within the quantum regime and provide energy storage and power delivery.
-We investigate the interplay between the battery and charger Hamiltonians, with a particular focus on minimizing the battery’s influence during the charging dynamics.
-To achieve this, we introduce a control parameter that allows us to suppress the battery’s contribution during the charging dynamics.
-We explore various configurations, including a non-interacting many-body battery with an interacting many-body charger, an interacting battery with a non-interacting charger, and systems where both the battery and charger are interacting many-body systems.
-Our results reveal a notable enhancement in stored energy and charging power when the battery’s influence is suppressed, underscoring the pivotal role of the charger in optimizing performance.
-Remarkably, across all scenarios, we observe that the presence of the battery’s countereffect within the charger Hamiltonian consistently leads to improved storage characteristics, highlighting a fresh direction in designing efficient quantum batteries.
+In this work, we explore the interplay between the battery and charger Hamiltonians, focusing on controlling and minimizing the battery’s intrinsic influence during the charging process.
+To this end, we introduce a tunable parameter that allows partial suppression of the battery’s contribution, enabling a systematic study of its role in energy transfer.
+We examine several charging configurations: a non-interacting qubit battery driven by an interacting many-body charger, an interacting qubit battery energized by a non-interacting charger, and setups in which both the battery and the charger are interacting qubit chains.
+In all cases, the inclusion of a controllable counteraction, or anti-effect of the battery Hamiltonian, allows us to modulate the battery’s internal dynamics during charging.
+Our results demonstrate a significant enhancement in both stored energy and charging power when the battery’s influence is suppressed, emphasizing the critical role of the charger in optimizing performance.
+Notably, we find that incorporating the battery’s countereffect consistently improves storage characteristics across all configurations, suggesting a novel avenue for designing highly efficient quantum batteries.
 I Introduction
-A quantum battery is a device designed to harness quantum properties to improve the efficiency of energy storage and retrieval.
-Traditional batteries store energy in chemical reactions [1, 2, 3, 4, 5], while quantum batteries would utilize the principles of quantum mechanics to store and release energy in a more efficient and potentially higher-capacity manner [6, 7, 8, 9, 10, 11, 12, 13].
-The basic idea behind a quantum battery is that quantum systems can exist in superposition states, meaning they can be in a combination of multiple states simultaneously.
-This property could potentially allow quantum batteries to store energy in a more distributed and efficient way compared to traditional batteries.
-Additionally, quantum entanglement, where particles become correlated in such a way that the state of one particle determines the state of the other particle, might enable quantum batteries to further enhance their energy storage capacity.
-Quantum batteries have attracted significant attention in recent years [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21].
-The primary goal of these efforts is to demonstrate how quantum mechanical systems and operations can enhance the stored energy or charging power of a battery.
-A variety of approaches—ranging from single-qubit protocols to many-body systems—have been explored to showcase this quantum advantage [10, 19, 20, 22, 16, 8, 23].
-Notable models include the Dicke quantum battery, which achieves modest speedups due to collective effects [16, 17], and qubit-based schemes that demonstrate an $N$-fold power advantage, pointing to strong scalability [8].
-Moreover, coherent quantum charging protocols have shown superior performance compared to classical strategies, with significantly enhanced power and robustness against decoherence and loss [24].
-Recent studies have also underscored the importance of interaction topology: optimal performance in fermionic systems is observed when the interaction connectivity aligns with the interaction order [23].
-In addition, alternative charging mechanisms such as those based on the Jaynes–Cummings model reveal that non-Gaussian cavity states can minimize energy fluctuations and achieve near-perfect charging fidelity [25].
-These theoretical developments have been paralleled by experimental advances, with successful realizations using superconducting circuits and trapped ions [26, 27, 28, 29].
-Periodic modulation techniques have also been applied to spin systems for efficient charging [30], and the role of dissipation in open quantum systems has been investigated as a tool for enhancing performance [31, 32, 33, 34, 35, 36, 37, 38, 12].
-Notably, global charging methods—where all cells are charged simultaneously—have been shown to offer the strongest quantum advantage, with power scaling quadratically with the system size [21].
-There has been a lack of discussion on the combined effect of the battery and charger in the charging dynamics, particularly in the context of spin systems.
-Addressing this gap, our study explores the interplay between the battery and charger during the charging process, emphasizing their joint influence on energy storage and power output.
-Our objective is to understand how optimizing their interactions can enhance overall performance while overcoming inherent limitations in the battery’s contribution.
-In quantum spin systems, both the battery and the charger play a role in charging dynamics for storing energy and power.
-However, we observe that the battery’s ability to achieve maximum storage energy and power is often constrained by the charger, which diminishes the battery’s effect during the charging process.
-To address this, we propose introducing an additional term in the charger Hamiltonian to counteract this limitation.
-By incorporating the battery’s countereffect into the charger, we enhance both storage energy and power across all considered configurations.
-This approach shifts the primary responsibility for energy storage to the charger, allowing the battery to adopt a more passive role, thereby optimizing the charging process and maximizing the efficiency of quantum batteries.
-The manuscript is organized as follows: Section II introduces the setup involving the combined battery and charger system and defines the key physical quantities used in the analysis.
+A quantum battery is a device designed to exploit quantum mechanical principles to enhance the efficiency of energy storage and retrieval.
+Unlike traditional batteries, which rely on chemical reactions [liu2014comparative, chaoui2017aging, chen2012state, fleischer2013adaptive, rivera2017soc], quantum batteries utilize features such as superposition and entanglement to store and release energy more efficiently and with potentially greater capacity [alicki2013entanglement, hovhannisyan2013entanglement, binder2015quantacell, campaioli2017enhancing, le2018spin, ghosh2020enhancement, farina2019charger, PhysRevA.105.022628].
+The ability of quantum systems to exist in superposed states allows energy to be distributed in ways not possible classically, while entanglement may provide additional advantages in energy storage, although it alone does not guarantee improved performance [campaioli2018quantum, gyhm2024beneficial].
+Extensive research has focused on demonstrating how quantum effects can enhance both the stored energy and the charging power of these systems.
+Various protocols—from single-qubit setups to complex many-body systems—have been developed to highlight this quantum advantage [le2018spin, julia2020bounds, rossini2020quantum, rosa2020ultra, ferraro2018high, binder2015quantacell, francica2024quantum].
+Models such as the Dicke quantum battery show collective effects that yield modest speedups [ferraro2018high, andolina2019quantum], while qubit-based schemes exhibit power scaling linearly with the number of cells, indicating strong scalability [binder2015quantacell].
+Coherent charging methods outperform classical approaches, providing enhanced power and robustness against noise and decoherence [salvia2023quantum].
+Recent findings also highlight the importance of interaction topology, revealing that optimal charging in fermionic systems occurs when the connectivity matches the interaction order [francica2024quantum].
+Additionally, charging mechanisms rooted in the Jaynes–Cummings model show that non-Gaussian cavity states can reduce energy fluctuations, achieving near-perfect charging fidelity [rinaldi2024reliable].
+These theoretical advances are supported by experimental progress in superconducting circuits and trapped ion platforms [quach2022superabsorption, hu2022optimal, dou2023superconducting, joshi2022experimental].
+Techniques such as periodic modulation of spin systems [mondal2022periodically] and controlled dissipation in open quantum systems have also been explored as means to enhance charging efficiency [ghosh2021fast, liu2024better, santos2021quantum, bai2020floquet, quach2020using, carrega2020dissipative, zhao2021quantum, barra2019dissipative, farina2019charger].
+Among these, global charging—where all battery units are charged simultaneously—has been found to deliver the strongest quantum advantage, with power scaling quadratically with system size [gyhm2022quantum].
+The investigation of many-body quantum systems remains central to understanding and optimizing quantum battery performance.
+Foundational studies have explored the XXZ spin chain with both local and non-local interactions [le2018spin], the Hubbard model, and bosonic and fermionic lattices [konar2022quantum].
+Spin-photon hybrid systems exhibit complex dynamical behaviors and rich interplay between components [ferraro2018high, crescente2020ultrafast, dou2022cavity, rodriguez2024optimal, erdman2024reinforcement, zhang2023enhanced, salvia2023quantum, hogan2024quench], while non-Hermitian models provide insight into energy transport and dissipation in open quantum systems [konar2024quantum].
+Comparisons with classical analogues demonstrate superior quantum retention and transfer efficiencies [andolina2019quantum], influenced by factors such as impurity-induced dimensional effects [PhysRevA.105.022628].
+Interestingly, environmental noise, typically viewed as detrimental, can under specific conditions accelerate the charging process [ghosh2021fast].
+A key element underpinning many quantum battery protocols is the utilization of quantum correlations, such as entanglement [hovhannisyan2013entanglement, campaioli2017enhancing, francica2017daemonic] and quantum coherence [gumberidze2019measurement, gumberidze2022pairwise, tirone2025quantum, catalano2024frustrating], which have been extensively investigated for their role in performance enhancement.
+Experimental realizations of these principles span platforms including superconducting circuits [hu2022optimal], semiconductor quantum dots [maillette2023experimental], and nuclear magnetic resonance setups [joshi2022experimental].
+The theoretical progress has been further extended to protocols for work extraction from partially characterized quantum energy sources, broadening the operational framework of quantum batteries [joshi2024maximal].
+Despite significant progress in quantum battery research, the coupled dynamics between the battery and charger, especially in spin-based quantum systems, have not been thoroughly explored.
+In this work, we address this gap by explicitly analyzing the interplay between the battery and charger Hamiltonians, showing how their interactions collectively determine the energy storage capacity and charging power.
+Our results reveal that although both the battery and charger are essential for the charging process, suppressive effects originating from the battery itself can limit its ability to maximize stored energy and power.
+As the charging Hamiltonian is activated, the battery naturally resists the incoming energy from the charger, which can limit the overall charging efficiency.
+To overcome this limitation, we introduce a modification to the self-contained battery–charger setup by incorporating the countereffect of the battery Hamiltonian, controlled by a tunable parameter $\lambda$.
+Experimentally, this countereffect can be implemented by dynamically shifting the energy levels of the battery spins when the battery is non-interacting, or by tuning the interaction strength when the battery is interacting, during the charging process.
+By including this back-action, the battery partially relinquishes its resistance to energy flow, allowing the charger to deliver energy more effectively.
+This strategy results in significant enhancements of both the stored energy and the charging power across a wide range of system configurations.
+Consequently, the battery assumes a more passive yet optimized role, enabling a more efficient and controllable charging protocol.
+Our framework thus provides a robust approach for improving quantum battery performance and offers a practical pathway toward high-efficiency energy storage at the quantum scale.
+The manuscript is organized as follows: Section II introduces the setup involving the combined battery and charger system, and defines the key physical quantities used in the analysis.
 Section III presents the results of energy storage and power calculations.
-Specifically, Subsection III.1 examines the scenario where the battery is non-interacting and the charger is interacting, while Subsection III.2 addresses the case of an interacting battery coupled to a non-interacting charger.
-Subsection III.3 explores the dynamics when both the battery and the charger are interacting.
-Finally, Section IV summarizes the main findings and concludes the manuscript.
+Specifically, section III.1 examines the scenario where the battery is non-interacting and the charger is interacting, while section III.2 addresses the case of an interacting battery coupled to a non-interacting charger.
+Section III.3 explores the dynamics when both the battery and the charger are interacting.
+Finally, section IV summarizes the main findings, and concludes the manuscript.
 II Set-up
-A quantum battery typically comprises a finite number of quantum systems, such as spins, that can store and transfer energy at the quantum level.
-The charger, also a quantum system composed of spins, is designed to inject energy into the battery.
-The dynamics of these spin-based systems are described by their respective Hamiltonians.
-The battery Hamiltonian, $\hat{H}_{B}$, describes the internal energy levels and/or interactions within the battery, whereas the charger Hamiltonian, $\hat{H}_{C}^{e}$, encapsulates the dynamics and properties of the charger.
-To initiate the charging process, the charger is connected to the battery, allowing them to interact.
-This interaction results in the formulation of the total charging Hamiltonian, which combines the contributions from both the battery and the charger.
-It is expressed as the sum of the battery Hamiltonian, $\hat{H}_{B}$, and the charger Hamiltonian, $\hat{H}_{C}^{e}$, capturing the dynamics of energy exchange between the two systems.
-Mathematically, it is expressed as
-$\hat{H}=\hat{H}_{B}+\hat{H}_{C}^{e}.$ (1)
-To extend the formulation, we treat the charger Hamiltonian in a unique way by considering both interacting and non-interacting forms.
-In each case, we incorporate the countereffect of the battery into the charger Hamiltonian to account for the mutual influence between the battery and charger during the charging dynamics.
-The strength of this countereffect is controlled by a parameter, $\lambda$, referred to as the battery’s countereffect parameter.
-This modified Hamiltonian is termed the effective charger Hamiltonian, as it accounts for both the intrinsic dynamics of the charger and the influence of the battery.
-Mathematically, it is given as:
-$\hat{H}_{C}^{e}=\hat{H}_{C}-\lambda\hat{H}_{B}.$ (2)
-This effective charger Hamiltonian is introduced during the charging process, where it operates alongside the battery Hamiltonian.
-By incorporating the battery’s countereffect into the charger dynamics, the total Hamiltonian governing the charging process is updated.
-Substituting the effective charger Hamiltonian into the total Hamiltonian [Eq.
-1], we obtain:
-$\hat{H}=\hat{H}_{B}+\hat{H}_{C}-\lambda\hat{H}_{B}=\hat{H}_{C}+(1-\lambda)\hat{H}_{B},$ (3)
-123NN-145
-Figure 1: Illustration of all-to-all interaction of spins.
-Physically interesting range of $\lambda$ lies between $[0,1]$.
-When $\lambda=0$, the battery fully contributes to energy storage and power during the charging process.
-Conversely, when $\lambda=1$, the battery’s influence is completely nullified during charging.
-For intermediate values of $\lambda$, the battery’s effect is progressively reduced, with the extent of this reduction determined by the strength of $\lambda$.
+We consider a quantum battery composed of a finite ensemble of qubits that can store and exchange energy through coherent quantum dynamics.
+Depending on the configuration, the qubits may either interact with one another or behave as noninteracting units, allowing us to explore how inter-qubit correlations influence the charging performance.
+The same ensemble of qubits also serves as the charger, meaning that both energy storage and charging processes are realized within a single physical system.
+This self-contained design enables internal charging driven by coherent many-body dynamics rather than by coupling to an external system.
+The overall evolution is governed by two Hamiltonians acting on the same set of qubits: the battery Hamiltonian, $\hat{H}_{B}$, which defines the intrinsic energy structure and correlations of the qubits, and the charger Hamiltonian, $\hat{H}_{C}$, which injects energy into the system during the charging interval.
+This unified framework provides a versatile platform to investigate how interactions, coherence, and quantum correlations determine the storage energy, charging power, and efficiency of the quantum battery.
+At the beginning of the protocol, the ensemble of qubits are initialized in an eigenstate, typically the ground state of the Ising battery Hamiltonian $\hat{H}_{B}$.
+At $t=0$, the charging Hamiltonian $\hat{H}_{C}$ is switched on, and the system evolves under the total Hamiltonian for a finite duration $t_{\mathrm{on}}$.
+During this charging interval, $\hat{H}_{C}$ drives the system out of equilibrium, enabling energy transfer into the battery degrees of freedom.
+After $t_{\mathrm{on}}$, $\hat{H}_{C}$ is switched off, and the system either continues to evolve under $\hat{H}_{B}$ alone or is measured to determine the stored energy.
+The instantaneous battery energy $\langle\hat{H}_{B}\rangle_{t}$, the stored energy $\Delta E(t)=\langle\hat{H}_{B}\rangle_{t}-\langle\hat{H}_{B}\rangle_{0}$, and the charging power are used to characterize the performance of the battery.
+The charging process is efficient because the two Hamiltonians do not commute, i.e., $\hat{H}_{B},\hat{H}_{C}]\neq 0$.
+This non-commutativity opens a microscopic channel for energy redistribution: the terms in $\hat{H}_{C}$ generate flip-flop excitations that propagate through the spin chain, transferring population to higher-energy eigenstates of $\hat{H}_{B}$.
+Once $\hat{H}_{C}$ is turned off, the injected energy remains stored within the ensemble of qubits, where it can either delocalize or remain localized, depending on the interaction strength and correlation structure of the system.
+When the charging Hamiltonian is activated, the battery Hamiltonian naturally resists the inflow of energy from the charger.
+To model a controlled suppression of the battery’s active response during charging, we introduce an externally applied anti-effect term.
+Specifically, when the charging Hamiltonian $\hat{H}_{C}$ is on, an additional term $\lambda\hat{H}_{B}$ is included in the total Hamiltonian.
+This term does not originate from the intrinsic battery dynamics but represents an external control that partially counteracts the internal battery Hamiltonian during the charging process.
+Experimentally, this can be implemented by shifting the energy levels of the non-interacting battery spins, or by tuning the interaction strength in an interacting battery.
+Physically, this mechanism mimics a feedback-like process that limits the battery’s response to the charger, effectively constraining it to absorb energy less actively.
+The strength of this counteraction is governed by the dimensionless parameter $\lambda$, with larger values corresponding to stronger suppression of the battery’s internal contribution.
+Consequently, the time-dependent Hamiltonian describing the charging pulse protocol takes the form:
+$\hat{H}(t)=\begin{cases}\hat{H}_{B},&t<0,\\[4.0pt] (1-\lambda)\hat{H}_{B}+\hat{H}_{C},&0\leq t\leq t_{\mathrm{on}},\\[4.0pt] \hat{H}_{B},&t>t_{\mathrm{on}}.\end{cases}$ (1)
+The parameter $\lambda\in[0,1]$ thus interpolates between the ideal charging limit ($\lambda=0$) and the fully suppressed limit ($\lambda=1$).
+During the charging pulse, $\hat{H}_{C}$ creates excitations that drive the system out of equilibrium, while the reduced Ising contribution $(1-\lambda)\hat{H}_{B}$ determines how efficiently those excitations are converted into stored energy.
+The externally applied countereffect term $\lambda\hat{H}_{B}$ can be implemented experimentally through different approaches, depending on whether the battery Hamiltonian is noninteracting or interacting.
+For a noninteracting battery defined by a local field $h$ and for an interacting battery characterized by coupling strength $J$, the countereffect can be realized in both cases by effectively reducing the active contribution of $\hat{H}_{B}$ during charging.
+Below, we discuss these processes, specifying where the anti-effect is applicable and how it can be implemented.
+Implementation of the noninteracting Battery Countereffect
+Since the battery Hamiltonian for the non-interacting qubits is simply a sum of local terms, i.e.
+$\hat{H}_{B}=h_{B}\sum_{j=1}^{N}\sigma_{j}^{z},$ the countereffect can be applied locally on each qubit during the charging process.
+This is achieved by introducing an additional energy shift or detuning along the $z$-axis for each battery qubit, effectively realizing the counterterm $\hat{H}_{B}^{\prime}=-\lambda\hat{H}_{B}=-\lambda h\sum_{j=1}^{N}\sigma_{j}^{z},$ where $\lambda\in[0,1]$ controls the strength of the countereffect.
+Physically, this corresponds to adding a magnetic field or qubit detuning $\delta_{j}=-\lambda h$ on each battery qubit.
+The following provides a brief overview of experimental platforms capable of realizing the externally applied anti-effect.
+  • Superconducting qubits: Per-qubit flux tuning can shift the qubit transition frequency by $\delta_{j}=-\lambda h$, providing a direct and controllable realization of the countereffect [paauw2009tuning].
+  • Trapped ions: AC-Stark shifts on individual ions can generate local $z$-fields of tunable magnitude and sign, implementing the desired energy shifts [staanum2002quantum].
+  • Cold atoms or Rydberg arrays: Site-selective optical potentials or Raman beams can shift the on-site energies of atoms representing the battery qubits, realizing the countereffect locally [griffin2005spatially].
+This direct Hamiltonian engineering approach allows for continuous tuning of the battery’s contribution to the charging dynamics, from fully active ($\lambda=0$) to completely suppressed ($\lambda=1$), providing a versatile and experimentally feasible method to explore different charging regimes.
+Interaction Control Implementation of the Countereffect
+An alternative and experimentally accessible method to implement the battery’s counter‑effect is via direct control of the qubit–qubit interactions within the battery subsystem.
+Many quantum platforms such as superconducting qubits, trapped ions, and Rydberg atom arrays—offer the capability to dynamically tune the effective coupling strength between qubits on timescales comparable to the charging interval.
+Concretely, suppose the battery qubits are coupled via an Ising-type Hamiltonian with interaction strength $J$.
+We incorporate the counter‑effect parameter $\lambda\in[0,1]$ by effectively suppressing a fraction $\lambda$ of that coupling: $J\;\longrightarrow\;(1-\lambda)\,J\,.$ Here $\lambda=0$ corresponds to the original full interaction (no suppression), whereas $\lambda=1$ corresponds to complete cancellation of the coupling.
+Physically, this modulation reduces the battery’s internal interaction, thereby diminishing its intrinsic energy‑redistribution and cooperative charging effects during the dynamics.
+The platforms where this externally applied anti-effect can be implemented are briefly outlined below.
+  • Superconducting qubits.
+Tunable couplers allow in situ variation of the coupling strength between qubits.
+For example, Harris et al.
+demonstrated a sign‑ and magnitude‑tunable coupler for superconducting flux qubits, enabling $J$ to be tuned from roughly +45+45 mK to 55-55mK (i.e., coupling can be turned off or even inverted) [harris2007sign].
+More recently, Li et al.
+implemented a flux‑controlled tunable coupler that can fully turn off adjacent qubit coupling and thus implement $(1-\lambda)J$ in practice [li2020tunable].
+  • Trapped ions.
+Laser‑induced spin–spin interactions mediated via collective motional modes can be tuned in strength and range.
+For instance, entangling interactions between two trapped ions were shown to be tunable by controlling the local potentials and coupling strength [wilson2014tunable].
+More broadly, programmable many‑body interactions in ion arrays have been demonstrated [katz2023programmable].
+  • Rydberg atom arrays.
+Modulation techniques (such as Floquet driving) allow switching between blockade and anti‑blockade regimes, effectively changing interaction strengths dynamically [zhao2023floquet].
+Although direct nearest‑neighbour $J$-tuning in this context is less often reported, the conceptual tools are available for similar interaction‑control protocols.
+Because this method directly manipulates the existing interaction Hamiltonian of the battery, without introducing additional couplings or measurement‑based feedback loops, it is conceptually straightforward and introduces minimal extra back‑action.
+Through tuning $\lambda$ from 0 to 1, one can experimentally explore the full spectrum of charging behaviour—ranging from strongly correlated, cooperative storage dynamics to suppressed or nearly non‑interacting charging regimes.
 To analyze the counter effect of the battery during the charging process, we calculate two key metrics: storage energy and power.
@@
 $\Delta E=\mathrm{Tr}[\hat{\rho}(t)\hat{H}_{B}]-\mathrm{Tr}[\hat{\rho}(0)\hat{H}_{B}],$ (2)
-where $\hat{\rho}(0)=|\psi_{0}\rangle\langle\psi_{0}|$, and $|\psi_{0}\rangle$ is the ground state of the battery Hamiltonian.
-The time evolution of the system’s state is given as: |$\ket{\psi(t)}=\hat{U}(t)\ket{\psi_{0}}$, where $\hat{U}(t)=e^{-i\hat{H}t}$ is the time-evolution operator.
+[…]
 The storage power quantifies the rate at which energy is transferred into the system during the charging process and is given by:
@@
 III result
+Several studies on many-body quantum batteries have have been done [le2018spin, ghosh2020enhancement, PhysRevA.105.022628], with particular emphasis on interacting spin-chain batteries.
+These works demonstrate that spin–spin interactions, disorder, and higher spin degrees of freedom play a crucial role in shaping the charging dynamics and improving power performance.
+Our scenario is slight different, as we focus on the role of the battery in the charging dynamics.
+In particular, we investigate how the intrinsic properties of the battery influence the overall charging process.
 To evaluate the performance of the quantum battery, we compute the stored energy and charging power across three distinct scenarios.
@@
 $\hat{H}_{B}=h\sum_{n=1}^{2^{N}}E_{n}|E_{n}\rangle\langle E_{n}|,$ (5)
-where $E_{n}$ are the eigenvalues, ranging from $-N$ (ground state energy) to $N$ (highest energy level).
+where $E_{n}$ are the eigenvalues, ranging from N-N (ground state energy) to NN (highest energy level).
 The total number of energy levels is $2^{N}$, with the ground state and the highest energy state being non-degenerate.
 The remaining energy levels exhibit degeneracy, given by $^{N}C_{l}$, where $l=1,2,\dots,N-1$, corresponding to intermediate energy levels.
-The goal of charging the battery is to transition it from the ground energy state |$\ket{E_{1}}$ with energy $-Nh$ to the highest energy state |$\ket{E_{2^{N}}}$ with energy $Nh$.
-The maximum storage energy $\Delta E_{\rm max}$ is defined as:
+123NN-145
+Figure 1: Illustration of an all-to-all interaction among spins, where every spin is coupled to all the others simultaneously.
+Such a configuration highlights the collective nature of spin dynamics compared to local or nearest-neighbor interactions.
+[…]
 […]
@@
 The maximum storage energy of the non-interacting battery, as described by Eq. (4), is given by $2Nh$.
-Figure 2: Non-interacting battery with ATA interacting Ising spin charger.
-(a) $\Delta E$ and (b) $P$ vs.
-$t$ for different values of $\lambda$ with fixed system size $N=10$.
-(c1) and (c2) show $\Delta E$ vs.
-$t$ for systems with (c1) odd and (c2) even $N$ with fixed $\lambda=1$.
-(d) $P$ vs.
-$t$ for various $N$.
-The parameters used are $J=h=1$, with periodic boundary conditions applied.
-Figure 3: Maximum storage (a) energy and (b) power as functions of the battery’s countereffect strength $\lambda$, for two interaction scenarios and their special cases, with system size fixed at $N=10$.
-Maximum storage (c) energy and (d) power as functions of system size $N$, with the battery counter strength fixed at $\lambda=1$.
-In all cases, the parameters used are $J=h=1$, and $\gamma=0.5$.
-To examine the influence of the battery’s countereffect on its storage energy and power, we consider an all-to-all (ATA) interacting Ising spin system as an interacting charger, as illustrated in Fig.
+For the charging the noniteracting battery, during the chaging process we switch one the all-to-all (ATA) interacting Ising spin system as an interacting charger, as illustrated in Fig.
 1.
@@
 We assume periodic boundary conditions, i.e., $\hat{\sigma}_{N+k}=\hat{\sigma}_{k}$, to preserve translational symmetry and eliminate edge effects.
-To avoid redundancy under periodic boundary conditions, we fix $\mathcal{K}=\frac{N-1}{2}$ for odd $N$ and $\mathcal{K}=\frac{N}{2}$ for even $N$, ensuring that each spin pair is included only once.
+To avoid redundancy under periodic boundary conditions, we fix $\mathcal{K}=\frac{N-1}{2}$ for odd NN and $\mathcal{K}=\frac{N}{2}$ for even $N$, ensuring that each spin pair is included only once.
+In this scenario, the battery is non-interacting while the charger is interacting.
+To incorporate the countereffect of the battery during the charging process, we introduce an additional energy shift in the battery Hamiltonian of the form $(1-\lambda)\hat{H}_{B}$.
+This externally applied term partially suppresses the active response of the battery spins, effectively reducing their resistance to energy inflow from the charger.
+Experimentally, for a non-interacting battery, this can be realized by applying a controlled detuning or local energy shift on each battery qubit, corresponding to the strength of $(1-\lambda)\hat{H}_{B}$.
+Physically, this procedure acts as a feedback-like mechanism, allowing the charger to inject energy more efficiently while the battery assumes a more passive but optimized role.
+The dimensionless parameter $\lambda\in[0,1]$ governs the strength of this countereffect, with larger $\lambda$ leading to stronger suppression of the battery’s internal contribution.
+Consequently, the total time-dependent Hamiltonian governing the charging dynamics can be expressed as $\hat{H}(t)=(1-\lambda)\hat{H}_{B}+\hat{H}_{C},$ where $\hat{H}_{C}$ represents the interacting charger Hamiltonian driving the energy flow.
 We calculate the storage energy of the non-interacting battery Hamiltonian with the ATA connected Ising spin system as the charger Hamiltonian for different values of $\lambda$ ranging from $[0,1]$.
@@
 2(b)].
+Figure 2: Non-interacting battery charged by an ATA interacting Ising spin system.
+(a) Stored energy $\Delta E$ and (b) charging power $P$ as functions of time $t$ for different values of the countereffect strength $\lambda$, with system size fixed at $N=10$.
+(c1) and (c2) show the time evolution of the stored energy $\Delta E$ for systems with odd and even numbers of spins NN, respectively, with the countereffect strength fixed at $\lambda=1$.
+(d) Charging power $P$ as a function of time $t$ for different system sizes NN.
+All simulations are performed with interaction strength $J=1$, external field $h=1$, and periodic boundary conditions.
 The storage energy is analyzed as a function of the system size, with the parameter $\lambda$ fixed at $1$, revealing an odd-even effect in its behaviour.
@@
 2($c_{2}$)].
+This observed odd–even effect arises from the interplay of spin-spin correlations and pairing in finite-size spin systems.
+For even-numbered batteries, all spins can optimally align or pair during charging, resulting in higher stored energy.
+For odd-numbered batteries, an unpaired spin leads to slight frustration, reducing the achievable energy.
+Consequently, the maximum storage energy and charging power exhibit a systematic dependence on whether the battery contains an odd or even number of spins, reflecting the underlying quantum correlations.
+These Refs.
+[politi2009even, herviou2023even] discussed the odd-evn effect in the differnt context.
 The storage power of the battery is analyzed as a function of the system size with a fixed value of $\lambda=1$.
@@
 This indicates that the nature of the spin-spin interaction affects the storage capacity.
+Figure 3: (a) Maximum stored energy and (b) maximum power as functions of the battery’s countereffect strength $\lambda$ for a non-interacting battery charged by four types of chargers: Ising ATA, Ising NN, XY ATA, and XY NN, with system size $N=10$.
+(c) Maximum stored energy and (d) maximum power as functions of system size NN for the same four charger types, with the countereffect strength fixed at $\lambda=1$.
+The parameters are $J=1$, $h=1$, and $\gamma=0.5$.
 Furthermore, we investigate how the system size $N$ influences the maximum storage energy and power, keeping the countereffect parameter fixed at $\lambda=1$.
@@
 A similar odd-even behavior appears in the XY ATA interacting charger.
-For odd system sizes, the storage energy remains below half of the maximum value ($<hN$), while for even system sizes, it exceeds half the maximum ($>hN$).
+For odd system sizes, the storage energy remains below half of the maximum value (<hN<hN), while for even system sizes, it exceeds half the maximum (>hN>hN).
 In contrast, the XY NN configuration yields a storage energy that remains below $hN$ for all system sizes [Fig.
@@
 III.2 Interacting battery and non-interacting charger Hamiltonian
-Figure 4: (a) $\Delta E$ and (b) $P$ vs.
-$t$ for NN Ising interacting Hamiltonian as a battery and non-interacting Hamiltonian as a charger.
-(c) $\Delta E$ and (d) $P$ vs.
-$t$ for NN interacting XY spin system as a battery and non-interacting Hamiltonian as a charger.
-Parameterse are: $N=12$, $J=1$, $h=1$, and $\gamma=0.5$.
-Periodic boundary conditions are considered.
+Figure 4: (a) Stored energy $\Delta E$ and (b) charging power $P$ as functions of time $t$ for a battery described by a NN Ising Hamiltonian with a non-interacting charger.
+(c) Stored energy $\Delta E$ and (d) charging power $P$ for a battery described by a NN XY Hamiltonian with the same non-interacting charger.
+The system parameters are $N=12$, $J=1$, $h=1$, and $\gamma=0.5$, with periodic boundary conditions.
 To generalize our observation that the countereffect of a battery in the charging dynamics enhances storage energy and power, we now consider a reversed setup: an interacting battery coupled to a non-interacting charger.
@@
 (4), i.e., $\hat{H}_{C}=h\sum_{j=1}^{N}\hat{\sigma}_{j}^{z}$.
-We analyze the storage energy and power of the battery in two scenarios: one where the battery’s influence on the charger is present ($\lambda=0$) and another where it is absent ($\lambda=1$).
-The results demonstrate that, for both Ising and XY interacting batteries, storage energy and power are notably higher when $\lambda=1$, underscoring the beneficial role of the battery’s countereffect in the charging dynamics.
+Within this model, the quantum battery is modeled as an interacting chain of qubits whose internal spin–spin couplings facilitate collective energy storage and the buildup of correlations during the charging process.
+When the charging protocol is initiated at time $t_{\mathrm{on}}$, a noninteracting charger—realized as a uniform external field—injects energy coherently into all qubits without generating additional correlations.
+To account for the feedback-like countereffect during charging, the internal interaction strength of the battery is simultaneously modulated as $J\rightarrow(1-\lambda)J$, where $\lambda$ quantifies the degree of suppression of the battery’s intrinsic contribution.
+This configuration, representing an interacting battery driven by a noninteracting charger, provides a clear framework to explore how intrinsic interactions influence the redistribution of injected energy, thereby enhancing storage capacity, charging power, and the overall robustness of the quantum battery.
+We analyze the storage energy and power of the battery in two scenarios: one where the battery’s countereffec on the charger is absent ($\lambda=0$) and another where it is present ($\lambda=1$).
+The results demonstrate that, for both Ising and XY interacting batteries, storage energy and power are notably higher when battery counter effect is maximal in the charger ($\lambda=1$), underscoring the beneficial role of the battery’s countereffect in the charging dynamics.
 Additionally, the Ising model consistently yields superior performance over the XY model in both metrics, as shown in Fig.
 4(a–d).
+We investigate how the storage energy and power depend on the interaction strength of the battery.
+In this analysis, the countereffect of the battery on the charger is not included, so as to clearly isolate the influence of interactions on storage energy and power.
+Our results show that the storage energy increases with $J$, reaches a maximum at a particular time, and then decreases, while the time required to attain the maximum also decreases as the interaction strength $J$ increases (Fig.
+5(a)).
+The inset of Fig.
+5(a) highlights how the maximum storage energy varies with $J$: it grows as $J$ approaches $1$, reaches its peak at $J=1$, and then decreases beyond this point.
+Notably, the critical value $J=1$ coincides with the charger strength $h=1$, suggesting that optimal storage energy is achieved when the battery and charger strengths are equal.
+The power follows a similar trend: it grows with time, reaches a maximum, and then decays for all values of $J$.
+Moreover, the time to reach the maximum power decreases as $J$ increases (Fig.
+5(b)).
+The maximum storage power exhibits a logarithmic dependence on $J$, approximately following $P_{\rm max}\approx 17.91\log_{10}(J)+16.86$ (inset of Fig.
+5(b)).
+Similar observations were reported in Ref.
+[ghosh2020enhancement], which demonstrated that both ordered and disordered interaction strengths significantly enhance power extraction.
+Additionally, Ref.
+[PhysRevA.105.022628] showed that the average power output of a quantum battery based on an interacting spin system, charged via a local magnetic field, increases with the spin quantum number, highlighting a dimensional advantage in quantum batteries.
+Figure 5: Storage energy $\Delta E$ and power $P$ of the NN Ising spin system as a quantum battery, with non-interacting spins as the charger, shown as a function of time for different interaction strengths $J$ (fixing $\lambda=0$).
+The inset shows the maximum storage energy and power as $J$ is varied.
+The maximum power scales approximately logarithmically with interaction strengh , following $P_{\rm max}\approx 17.91\log_{10}(J)+16.86$.
 III.3 Interacting battery and interacting charger Hamiltonian
-We consider the scenario where both the battery and charger are governed by interacting Hamiltonians.
-First, we analyze the case where the battery is modeled by an Ising spin system with NN interactions [Eq.
-(10)], and the charger is described by an NN XY-interacting Hamiltonian [Eq.
+Figure 6: (a) Stored energy $\Delta E$ and (b) charging power $P$ as functions of time $t$ for a battery described by a NN Ising spin system and a charger described by a NN XY spin system.
+(c) Stored energy $\Delta E$ and (d) charging power $P$ for the case where the battery is a NN XY spin system and the charger is a NN Ising spin system.
+The calculations use $N=12$, $J=1$, and $\gamma=0.5$ with periodic boundary conditions.
+We consider the scenario in which both the battery and the charger are interacting spin systems.
+In the first case, the battery is modeled as an Ising spin chain with NN interactions [Eq.
+(8)], while the charger is implemented as a nearest-neighbor XY-interacting Hamiltonian [Eq.
 (10)].
+In the reversed configuration, the roles are swapped: the battery is described by a NN XY-interacting chain, and the charger is realized as an NN Ising spin system.
+In this configuration, the system is initially prepared in the ground state of an Ising-type Hamiltonian, representing an interacting quantum battery characterized by spin–spin couplings that enable collective energy storage.
+At the onset of charging ($t=t_{\mathrm{on}}$), the interaction profile of the system is suddenly modified by switching on an XY-type Hamiltonian, which now acts as an effective charger.
+This sudden quench introduces flip-flop interactions that coherently inject energy into the spin system, promoting excitation delocalization and the build-up of correlations among the qubits.
+The two Hamiltonians—Ising and XY—thus represent distinct dynamical regimes of the same physical platform, allowing the same spin chain to serve sequentially as both the battery and the charger.
+To investigate the role of the countereffect, the intrinsic Ising interaction strength is simultaneously reduced as $J\rightarrow(1-\lambda)J$ during charging, where $\lambda$ controls the degree to which internal correlations are suppressed.
+This setup provides a unified framework to explore how the interplay between Ising and XY interactions governs energy absorption, redistribution, and storage efficiency, and how the externally applied countereffect modifies these cooperative charging dynamics.
 We calculate the storage energy and power in two situations: when the battery’s influence is present in the charging dynamics ($\lambda=0$) and when it is absent ($\lambda=1$).
 Our results indicate that the presence of the counter effect of battery during charging leads to an advantage in both the stored energy and the power [Fig.
-5(a,b)].
+6(a,b)].
 Next, we examine the reverse configuration: the battery is now represented by an NN XY-interacting spin system [Eq.
@@
 In this case, we again compute the storage energy and power, and we find that when the battery’s influence is absent in the charging dynamics ($\lambda=1$), the system exhibits improved performance in terms of both energy storage and power [Fig.
-5(c,d)].
+6(c,d)].
 In nutshell, across all scenarios considered, the presence of the battery’s counter effect in the charging dynamics consistently provides an advantage in both energy storage and power.
 This advantage is consistent for being independent of the interaction and noniteration of the battery and charger Hamiltonians.
-Figure 5: (a) $\Delta E$ and (b) $P$ as a function of time $t$ for a NN interacting Ising spin system as the battery and NN interacting XY spin system as the charger Hamiltonian.
-(c) $\Delta E$ and (d) $P$ vs.
-$t$ for a NN interacting XY spin system as the battery and NN interacting Ising spin system as the charger Hamiltonian.
-The system parameters are $N=12$, $J=1$, and $\gamma=0.5$.
-Periodic boundary conditions are considered.
 IV Conclusion
-In this work, we explored the role of the battery’s countereffect in the charging dynamics of quantum batteries composed of spin systems.
-By analyzing various configurations involving both non-interacting and interacting Hamiltonians for the battery and charger, we demonstrated how their interplay significantly impacts storage energy and power.
-We began by examining a non-interacting battery coupled to an interacting charger modeled by an ATA Ising spin system.
-Our results revealed that incorporating the battery’s countereffect into the dynamics ($\lambda=1$) leads to optimal performance, with the storage energy reaching its theoretical maximum $2Nh$ and the power also peaking.
-Notably, an odd-even parity effect emerged in energy storage for increasing system sizes, while power remained parity-independent.
-To further understand the influence of interaction types and ranges, we compared Ising and XY charger models with both ATA and NN interactions.
-Across all these cases, the optimum storage energy and power consistently occurred at $\lambda=1$, emphasizing the constructive role of the countereffect in enhancing energy transfer.
-Both quantities exhibited linear scaling with system size, confirming the scalability of our protocol.
-Extending the analysis to interacting batteries with non-interacting chargers, where the battery is governed by either Ising or XY NN interactions, we observed that the inclusion of the countereffect continued to offer superior energy and power outcomes.
-This advantage persisted even in the most general scenario where both the battery and charger are interacting systems, reinforcing the robustness of our findings across diverse interaction frameworks.
-Interestingly, in cases where the charger is modeled with ATA interactions, the charging power peaks beyond the physically meaningful range ($\lambda>1$).
-Although this shift lies outside the regime of practical relevance, it suggests the presence of underlying physical mechanisms that merit deeper investigation.
-A detailed discussion of this anomaly is provided in Appendix A.
-Altogether, our study establishes the critical importance of retaining the battery’s countereffect in the charger Hamiltonian for optimizing quantum battery performance.
-These insights offer valuable guidance for the design of efficient and scalable quantum energy storage devices and open new avenues for exploring interaction-driven enhancements in quantum thermodynamic protocols.
+This work presents a comprehensive study of charging dynamics in spin-based quantum batteries, focusing on the often-overlooked influence of the battery’s countereffect during the charging dynamics.
+By systematically analyzing configurations involving interacting and non-interacting Hamiltonians for both the battery and the charger, we reveal a consistent and compelling theme: retaining the battery’s countereffect significantly boosts charging performance.
+Our findings show that maximum storage energy and power are achieved when the countereffect is fully included ($\lambda=1$), regardless of whether the system features ATA or NN interactions, and whether the interactions are of the Ising or XY type.
+Particularly, in the case of a non-interacting battery with an ATA Ising charger, the stored energy reaches its upper bound $2Nh$, accompanied by a notable odd-even parity effect in energy scaling—though power remains unaffected by parity.
+Importantly, this enhancement remains robust when the roles are reversed, i.e., with an interacting battery and a non-interacting charger.
+Even in the most general case, where both battery and charger are interacting many-body systems, the countereffect continues to improve energy transfer and power, suggesting it as a universal design principle for efficient quantum batteries.
+In the case of a NN Ising interaction battery coupled to a non-interacting charger, the battery’s interaction strength enhances the maximum storage energy up to the point where the interaction strength matches that of the charger.
+However, the maximum power increases approximately logarithmically with the interaction strength.
+A curious feature emerges in some ATA models where peak power occurs at $\lambda>1$, outside the regime of physical interpretability.
+While not directly useful for device implementation, this behavior indicates nontrivial underlying dynamics, which could inspire future theoretical investigations (see Appendix A).
+Altogether, our results uncover a simple yet powerful insight: the battery’s back-action on the charger is not a hindrance but a resource.
+Incorporating this countereffect in the system design leads to substantial gains in energy and power storage, offering valuable direction for next-generation quantum energy storage technologies.
 Appendix A Charging dynamics beyond physical regime
-Figure 6: Storage power $P(t)$ as a function of time $t$ for ATA interacting (a) Ising and (b) $XY$ charger Hamiltonian, shown for various values of the countereffect parameter $\lambda$.
-The insets display the corresponding maximum storage power as a function of $\lambda$.
-Parameters used: $J=h=1$, $N=10$, and $\gamma=0.5$.
+Figure 7: Charging power $P(t)$ as a function of time $t$ for a non-interacting battery charged by an ATA interacting (a) Ising and (b) XY Hamiltonian.
+Results are shown for several values of the countereffect parameter $\lambda$.
+The insets display the corresponding maximum charging power as a function of $\lambda$.
+The simulations use $J=1$, $h=1$, system size $N=10$, and anisotropy parameter $\gamma=0.5$.
 In the main manuscript, we investigate the charging dynamics of quantum batteries within the physically acceptable range of the countereffect parameter, $\lambda\in[0,1]$.
@@
 In both cases, for all values of $\lambda$, the storage power exhibits a consistent trend: it initially increases, reaches a maximum value, and subsequently decreases after a certain time [Fig.
-6(a,b)].
+7(a,b)].
 The maximum storage power, $P_{\rm max}$, depends on $\lambda$ and increases with it, attaining its peak at $\lambda=1.6$ for the ATA interacting Ising charger and $\lambda=2.8$ for the ATA interacting XY charger.
 Beyond these points, the storage power decreases with further increases in $\lambda$ [inset of Fig.
-6(a,b)].
+7(a,b)].
 Remarkably, the highest $P_{\rm max}$ values occur outside the physically accepted regime ($\lambda>1$).